Can someone explain to me how to maximize by the simplex method? I get confused after the second step.
Here is the problem,
Maximize P=0.3x ++ 0.4y
Subject to: 4x + 8y equals or less than 1600
12x+8y equal or less than 1920
x equal to or greater than 0; y equal to or greater than 0
Please help me. I know how to do the first step of charting but the second step on confuses me? How do I decide which negative should be the pivot point?

Answer by stanbon(75887)

Here is the problem,
Maximize P=0.3x ++ 0.4y
Subject to:
4x + 8y equals or less than 1600
12x+8y equal or less than 1920
x equal to or greater than 0; y equal to or greater than 0
Please help me. I know how to do the first step of charting but the second step on confuses me? How do I decide which negative should be the pivot point?
------------------------------------------
Solve each of the inequalities for "y"
y <=(-1/2)x + 200
y <= (-3/2)x + 240
---------------------------
Plot those boundary lines and the intersecting solution regions.

--------------------------
Find the corners of the solution region:(0,0),(0,200),(40,180),(160.0)
----------------------
Check each of those number pairs in P=0.3x + 0.4y
to determine which pair give you the maximum "p" value.
(0,0) gives P = 0
(0,200) gives p = 0 + 0.4*200 = 80
(40,180) gives p = 0.3*40 + 0.4*180 = 84
(150,0) gives p = 0.3*150 + 0 = 45
--------------------------------------------
So (40,180) gives the maximum Profit.
===============================================
Cheers,
Stan H.